# Efficiently serving large multimedia models using EPD Disaggregation

### v1

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead. This step helps convert raw inputs into tokenized representations that inflate the token sequence for the prefill phase, negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput. We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources. Unlike current systems, which bundle encoding and prefill together, our disaggregation approach alleviates memory bottlenecks, mitigates synchronization delays, and supports flexible batching. Specifically, we employ a new caching mechanism for multimodal tokens, enabling asynchronous transfer of multimodal tokens and introduce an integrated module to find optimal config for EPD system and minimize resource usage while maximizing SLO-based performance metric. Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ lesser for encoding-stage GPUs), that supports upto 22$\times$ higher batch sizes, 10$\times$ more number of images/ request, 2.2$\times$ higher kv cache size. Further, it leads to significant improvements in end-to-end throughput (up to 57% better), and latency metrics (TTFT up to 71% lower), compared to systems that do not disaggregate. Our findings underscore the potential of EPD disaggregation to enable resource-efficient and high-performance multimodal inference at scale.
@@ 1 Introduction
In current systems, the encoding and prefill stages are typically aggregated into a single monolithic and synchronous step executed on the same set of GPUs. This approach assumes that combining encoding and prefill can simplify implementation, but in practice, it exacerbates inefficiencies such as memory bottlenecks, limited batching flexibility, and synchronization delays. Specifically, simultaneous loading of both the multimodal encoder and LLM onto the same GPUs restricts memory for higher-resolution inputs or larger batch sizes, while the sequential execution of encoding and prefill introduces delays for encoding-light requests that must wait for encoding-heavy ones in the same batch. As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, making it clear that current solutions are inadequate for LMM workloads.
To address these challenges, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encoding, prefill, and decode stages, assigning each with dedicated resources to operate independently. This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention. By separating the multimodal encoder from the LLM on distinct GPUs, the system significantly improves memory efficiency, allowing for larger batch sizes and higher-resolution inputs. Additionally, asynchronous token transfer between stages minimizes latency, ensuring smooth and efficient processing. As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
@@
  • We identify major challenges and performance bottlenecks due to which the performance of existing multimodal inference systems is suboptimal. Specifically, these stem from encoding and prefill being a single monolithic and synchronous step that results in poor performance owing to inefficient utilization of memory and computation resources.
  • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages. Our approach addresses key challenges of resource allocation, inter-stage communication, and performance optimization within this framework. Specifically, we propose allocation of a new cache to store multimodal tokens on encoding and prefill stage that allows asynchronous transfer of multimedia tokens in an efficient manner.
  • We formulate the search for the optimal configuration of the EPD disaggregated system as a black-box optimization problem. This approach is simple yet practical because, on the one hand, it allows for a flexible search space that easily incorporates realistic configurations, including various parallelization strategies, GPU allocation, batch size selection, and complex scheduling strategies. On the other hand, it is computationally efficient, as a simulation outputs performance metrics (e.g., goodput or E2ETP) without actually deploying the system.
  • We provide a detailed experimental analysis of our method with respect to baselines on various popular LMM models, namely MiniCPMv 2.6, InternVL2-8B, and InternVL2-26B. The results demonstrate the superiority of our method in terms various aspects, like memory (upto 15x lesser), facilitating upto 22x higger batch sizes, 10x more number of images/ request, 2.2x higher kv cache size. Further, our method achieves significantly improved performance in terms of SLO attainment (upto 90-100% more), TTFT latency (upto 71% lesser), and E2ETP (upto 57% better).
@@ 2 Related Work
Serving LLMs and LMMs, which often consist of billions of parameters, poses significant challenges due to slow inference speeds [12, 25]. Efficiently serving these models to billions of users while adhering to stringent Quality of Service (QoS) criteria has therefore become an active area of research [6]. This has led to the development of numerous production-grade systems, such as vLLM [10], TorchServe [17], NVIDIA Triton [13], SGLang [23], and HuggingFace Text Generation Inference (TGI) [18], that integrate ideas emerging from recent advancements in this field.
Throughput-Optimized Inference Systems: One prominent direction of research focuses on improving throughput, a critical system-level metric, by optimizing GPU utilization. For instance, Orca [22] proposed continuous batching, a technique that dynamically increases batch size without interrupting ongoing decoding, thereby enhancing GPU efficiency. vLLM [10] introduced paged KV cache management, which reduces memory fragmentation and frees up GPU memory, enabling larger batch sizes and ultimately boosting throughput. Similarly, SARATHI [2] introduced chunked prefill, which splits prefill requests into smaller chunks and piggybacks decoding requests alongside them to improve GPU utilization. However, these systems are designed for LLMs and do not account for the resource-intensive multimodal encoding step unique to LMMs, which precedes the prefill phase.
Latency-Optimized Inference Systems: Another line of work focuses on improving user-level metrics, such as request latency, for individual queries. For example, FastServe [19] employs iteration-level preemptive scheduling to reduce latency caused by head-of-line blocking in throughput-optimized systems. To gain finer control over TTFT and TPOT, approaches like SplitWise [14], DistServe [24], and DéjàVu [16] disaggregate the prefill and decode stages, mitigating interference between them. However, these methods are designed for LLMs and overlook the resource-intensive encoding step present in LMMs. Since encoding is coupled with prefill, these systems suffer from poor TTFT latency, particularly when the resolution and quantity of multimodal data are large.
We propose disaggregation of encoding from prefill, enabling fine-grained control over batching, scheduling, and parallelization strategies, thereby achieving better latency and throughput in LMMs.
@@ 3 Method
Figure 1: The inference pipeline of EPD Disaggregation system.
Figure 2: System architecture of the proposed EPD Disaggregated Inference. It consists of three stages. In each stage, data parallelism (DP) means multiple instances serve in parallel. Each instance encompasses pipeline-parallel (PP) workers and tensor-parallel (TP) workers. For the encoding stage, since intra-request parallel (IRP) is better than TP, we use IRP instead of TP workers. MM stands for Multi-Media.
3.1 LMM System
At a high level, the LMM system transforms multimodal data and a text prompt into a coherent textual output. The system receives three types of input: (1) a text prompt $i_{p}$, which provides the context for the task (e.g., “What is the content of the image?”), (2) multimodal data $i_{m}$, which includes images, videos, or other sensory inputs (e.g., one or more images of a cat), and (3) sampling data $i_{s}$, which consists of parameters such as the desired number of output tokens and temperature settings. The system processes these inputs to generate a text output $o$, which is the model’s response (e.g., “The image contains a cat.”). For simplicity, we focus on the visual modality in this paper, but the framework is extensible to other types of multimodal data as well.
@@ 3.2 EPD Disaggregation
Disaggregating an LMM system involves breaking the inference process into distinct stages, each with specific inputs and outputs. As shown in Figure  1, the system follows a multi-stage inference pipeline consisting of three primary components: encoding, prefill, and decode. Additionally, the transitions between these stages are managed through EP-migration and PD-migration, which handle the transfer of data between encoding-prefill and prefill-decode stages, respectively. In the following, we describe each stage and the transformations involved.
Encoding (E): The multimodal input $i_{m}$ is processed by the multimodal encoder (MME) $E$, which converts the input data into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs. These tokens represent a high-dimensional embedding of the input and serve as the foundation for the next stage of the pipeline.
$v_{t}^{e}=E(i_{m})$ (1)
EP-migration: Once the encoding stage $E$ completes processing, the generated multimodal tokens $v_{t}^{e}$ must be transferred to the prefill stage $P$. This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache.
$v_{t}^{p}=\psi_{EP}(v_{t}^{e})$ (2)
Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
Prefill (P): The multimodal tokens $v_{t}^{p}$ arrive at the Prefill stage $P$, along with the text prompt $i_{p}$. In this stage, the system initializes the Key-Value (KV) cache $kv_{1}^{p}$ and generates the first output token $o_{1}^{p}$. The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage.
$kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$ (3)
PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$. This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker.
$kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$ (4)
Decode (D): The Decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the previously generated token $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$. This autoregressive process continues until the desired output sequence is produced.
$kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$ (5)
Each token generation step depends on the previous token and the context stored in the KV cache, ensuring the system generates contextually accurate and coherent outputs.
@@ 3.3 System Design and Optimization
Figure 2 illustrates the architecture of our EPD Disaggregated Inference system. Each stage of the pipeline—Encoding, Prefill, and Decoding—features fully independent instances capable of running the corresponding stage. These instances are designed to operate in DP mode, meaning multiple instances per stage can process different requests concurrently, ensuring scalability and efficiency.
@@
In the Encoding Stage, instances/ workers only load the MME as the LLM is not required at this stage. Similarly, only the MM cache is necessary to be initialized. In the Prefill Stage, instances load the LLM to handle prefill operations, but the MME is no longer required. These instances utilize both the MM cache and the KV cache to efficiently manage multimedia and key-value data associated with the request. In the Decoding Stage, instances also load the LLM, focusing exclusively on decoding tasks. The MME is not required at this stage, and only the KV cache is utilized.
All cache transfers occur when an instance pulls a request from its respective queue, ensuring that data is moved only when necessary. Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on reducing latency (Section 3.3.2), managing resources effectively (Section 3.3.3), and ensuring smooth token transfers between stages (Section 3.3.1).
@@ 3.3.1 Asynchronous Token Transfer
To minimize latency during token transfers between stages, our system supports direct, asynchronous token transfers via high-bandwidth NVLink/ InfiniBand / NVSwitch channels. The transfer occurs asynchronously, allowing the system to continue processing new requests while the token transfer is in progress. Specifically, we maintain a multimodal cache (MM cache) on both encoding and prefill workers. When a request completes the encoding stage, its tokens are first stored in the encoding worker’s MM cache, freeing the encoding worker to serve subsequent requests immediately.
An independent asynchronous event loop continually monitors for completed encoding tasks. When a request’s tokens become available, this event loop initiates a direct transfer of the tokens to the prefill worker’s MM cache. Because the transfer is asynchronous, the system can process other tasks in parallel without stalling. Once the token transfer is confirmed, the corresponding MM cache entries at the encoding stage are cleared to reclaim memory for future requests.
To manage these cache regions efficiently, we introduce an MMBlockManager. For each incoming request, the MMBlockManager precomputes the number of cache blocks the request will require and allocates them accordingly. After the request transitions from the encoding to the prefill stage and the tokens are transferred, those blocks are reassigned or de-allocated, enabling flexible cache usage even under heavy workloads.
Figure 3: The illustration of Intra-Request Parallel.
@@ 3.3.2 Intra-Request Parallel (IRP)
Often, multimodal requests include multiple high-resolution images, each of which is partitioned into patches for high-resolution encoding. For modern LMMs, the encoder responsible for processing these patches can be large, making the encoding phase computation-bound. Hence, if a request is processed serially, the encoding phase can become a bottleneck, adding significant delay to the first token latency TTFT.
To address this, we introduce Intra-Request Parallel (IRP), which splits a single request’s image patches among multiple encoding workers operating in data-parallel mode. We are guided by the intuition that these patches are encoded independently and hence can be processed in a parallel and asynchronous manner. As depicted in Fig. 3, each encoding worker receives a distinct subset of patches, computes the corresponding token representations, and transfers them to the prefill stage independently and asynchronously. This means that patches assigned to one encoding worker can finish and begin transferring without waiting for patches processed by other workers.
Once all the patch-level tokens arrive at the prefill stage, they are aligned, projected, and merged to form the complete set of multimodal tokens for the request. The request then proceeds through the prefill stage as usual, with the corresponding LLM processing the aggregated multimodal and prompt tokens. By allowing concurrent encoding of patches across multiple workers, IRP effectively reduces end-to-end latency and improves the utilization of available encoding resources, especially for high-resolution inputs with large numbers of patches.

### v2

# Efficiently Serving Large Multimodal Models Using EPD Disaggregation

@@ Abstract
Large Multimodal Models (LMMs) extend Large Language Models (LLMs) by handling diverse inputs such as images, audio, and video, but at the cost of adding a multimodal encoding stage that increases both computational and memory overhead. This step negatively impacting key Service Level Objectives (SLOs) like time to first token (TTFT) and end-to-end throughput (E2ETP). We introduce Encode-Prefill-Decode (EPD) Disaggregation, a novel framework that separates the encoding, prefill, and decode stages onto dedicated resources. Unlike current systems, which bundle encoding and prefill together, our approach decouple these steps unlocking new opportunities and optimizations. These include a new mechanism to cache multimedia tokens for efficient transfer, a novel way to parallelize encoding load within a request, a module to find the optimal resource allocation for disaggregated serving, and a novel role switching method to handle changing workload characteristics. Experimental evaluations with popular LMMs show substantial gains in memory efficiency (up to 15$\times$ less utilization), batch sizes (up to 22$\times$ larger), 10$\times$ more images/request, and 2.2$\times$ larger KV caches. Further, it leads to significant improvements in latency metrics (TTFT up to 71% reduction) and end-to-end throughput (up to 57% reduction), compared to systems that do not disaggregate.
@@ 1 Introduction
Figure 1: Above: Aggregated system, where E and LLM are coupled and must run on the same GPU. Below: Disaggregated system, where E runs on the 1st GPUs and LLM runs on the rest of three GPUs. Multimodal encoder and LLM are denoted by E and LLM, respectively.
@@
Figure 2: Impact of disaggregation on batch size and number of images/ request for the MiniCPM-V 2.6 model
These challenges of LLM deployment present opportunities for optimization that current serving systems do not consider. Currently, the encode and prefill stages are aggregated into a single monolithic and synchronous step executed on the same set of GPUs. Sequential execution of encoding and prefilling introduces interference. This interference is demonstrated graphically in Figure 1. In the aggregated approach (Figure 1 top), the prefill step ($LLM^{4}$) interferes with the encode-heavy request ($E^{5}$). As a result, such aggregation leads to suboptimal resource utilization and poor SLO performance, revealing that current solutions are inadequate for LMM workloads. By disaggregating encode from prefill, the system may be optimized to reduce interferences for specific workloads. From another perspective, simultaneous loading of both the Multimodal encoder (MME) and LLM onto the same GPUs in aggregated setups restricts memory for processing multiple high resolution images or larger batch sizes. In preliminary investigations, by removing the LLM and associated KV cache, we observe a significant increase in maximum batch sizes and supported number of images/ request (Figure 2).
Thus, we propose Encode-Prefill-Decode (EPD) Disaggregation, a framework that decouples the encode, prefill, and decode stages, assigning each with dedicated resources to operate independently. This disaggregation enables customized batching, parallelization, and scheduling strategies for each stage, optimizing resource utilization while reducing contention. As a result, EPD disaggregation achieves better memory utilization, higher throughput, and improved compliance with critical SLOs like TTFT, TPOT, and E2ETP.
@@
  • We propose an efficient system for LMM inference that relies on the novel idea of disaggregating encoding and prefill stages. Our approach addresses key challenges of inter-stage communication, efficient parallelization, resource allocation, and performance optimization within this framework.
  • We propose intra-request parallelization (IPR) that shards a request into independent encoding jobs which can be executed in parallel, significantly reducing first token latency.
  • We formulate the resource allocation problem as an optimization over batch sizes, scheduling strategies, and parallelization approaches for each pipeline stage. Further, we provide a simple black box optimizer that uses workload samples to approximate the optimal config for a given workload.
  • We develop dynamic role switching capability that monitors the system for any bottlenecks and allows efficient reallocation of resources between E, P and D stages by efficiently switching ($\sim$0.7s) the role of an instance among these stages. This ensures system can respond to any changes in workload requirements during online serving.
  • We conduct evaluations on various popular LMM models, namely MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B using real-world and synthetic workloads. The results demonstrate the superiority of our method across various aspects. For instance, our approach requires up to 15x less memory, facilitating up to 22x larger batch sizes, 10x more images/request, and 2.2x larger KV caches. Further, our method significantly improves performance in terms of SLO attainment (up to 90–100% more) and TTFT latency (up to 71% reduction).
@@ 2 Related Work
In this section, we describe feasible approaches in the literature for multi-modality model serving and their limitations and then review the use of the disaggregation technique.
Multi-modality Model Serving. There are two ways to enable a multi-modality model serving system: (1) adopting and extending existing LLM serving systems, such as vLLM [9], SARATHI [2], and Orca [20]; and (2) leveraging the open-sourced code from recent works of improving multi-modality model inference. The first way fails to address the encoding bottleneck inherent in LMMs, limiting their applicability to multimodal workloads, particularly when handling high-volume and high-resolution multi-media data. In the second way, recent advancements such as KV cache eviction [10, 13] and compression [8] focus on specific features and limited scenarios. For instance, Inf-MLLM [13] enables efficient streaming inference for LMMs on a single GPU, targeting resource-constrained scenarios. While their implementations are feasible for serving LMMs, they fail to meet user SLOs in cloud serving scenarios. To the best of our knowledge, we are the first to propose an LMM serving system that leverages disaggregation and a series of integrated techniques to enable better resource allocation and improved SLOs.
Disaggregated Serving. Disaggregated serving has emerged as a promising technique in large model serving. By decoupling the prefill and decode stages, systems like SplitWise [14], DistServe [21], and DéjàVu [16] mitigate interference between these phases, enabling finer control over Time to First Token (TTFT) and Time Per Output Token (TPOT). However, these systems primarily target LLMs and overlook the encoding step required for LMMs, which is tightly coupled with prefill. Recent work, such as Mooncake [15] and PD-Serve [7], has extended disaggregation to include KV cache management and other system-level optimizations. Despite these advancements, they remain limited in their applicability to LMMs. Disaggregating the encoding phase from prefill could unlock new opportunities for optimizing LMM serving. For instance, when handling requests with long videos, encoding the frames could be executed in parallel, reducing head-of-line blocking and improving overall system efficiency. This approach would enable more flexible batching and scheduling strategies, ultimately achieving better latency and throughput for multimodal workloads.
@@ 3 Method
In this section, first, we describe the EPD disaggregation framework (Sec. 3.1), followed by a detailed explanation of the system design and optimizations (Sec. 3.2), that includes the asynchronous token transfer, intra-request parallel, optimized resource allocation, and dynamic role switching.
Figure 3: The inference pipeline of EPD Disaggregation.
@@ 3.1 EPD Disaggregation
As shown in Figure 3, disaggregating an LMM system involves breaking the inference process into multiple stages: encoding, prefill, and decode. Transitions between these stages—EP-migration and PD-migration—handle the transfer of data from encoding to prefill and from prefill to decode, respectively. Formally, we denote the input text prompt as $i_{p}$, multimodal data as $i_{m}$, and the text output as $o$, and provide details for each step and transition as follows:
Encoding (E): The multimodal input $i_{m}$ is processed by the MME $E$, converting it into multimodal tokens $v_{t}^{e}$ on encoding stage GPUs. These tokens form a high-dimensional multimodal embedding for the next pipeline stage. This process can be denoted as $v_{t}^{e}=E(i_{m})$.
EP-migration: Once the $E$ stage completes processing, the generated multimodal tokens $v_{t}^{e}$ are transferred to the prefill stage $P$. This transfer is managed by the EP-migration function $\psi_{EP}$, which handles the migration of tokens from the encoding worker’s cache to the prefill worker’s cache. This process is expressed as $v_{t}^{p}=\psi_{EP}(v_{t}^{e})$. Here, $v_{t}^{p}$ represents the tokens in the prefill stage after migration.
Prefill (P): The multimodal tokens $v_{t}^{p}$ arrive at the Prefill stage $P$, along with the text prompt $i_{p}$. In this stage, the system initializes the Key-Value (KV) cache $kv_{1}^{p}$ and generates the first output token $o_{1}^{p}$. The KV cache stores context and intermediate states, which are necessary for the subsequent decoding stage. The process is denoted by $kv_{1}^{p},o_{1}^{p}=P(v_{t},i_{p})$.
PD-migration: After the prefill stage generates the initial token $o_{1}^{p}$ and updates the KV cache to $kv_{1}^{p}$, the data must be transferred to the decode stage $D$. This transfer is managed by the PD-migration function $\psi_{PD}$, which ensures that the KV cache and the generated token are correctly handed off to the decode worker. This is denoted as $kv_{1}^{d},o_{1}^{d}=\psi_{PD}(kv_{1}^{p},o_{1}^{p})$.
Decode (D): The decode phase $D$ operates iteratively, using the KV cache $kv_{t}^{d}$ and the token generated in the previous iteration $o_{t}^{d}$ to generate the next token $o_{t+1}^{d}$ and update the KV cache to $kv_{t+1}^{d}$. This autoregressive process continues until the desired output sequence is generated, as denoted by $kv_{t+1}^{d},o_{t+1}^{d}=D(kv_{t}^{d},o_{t}^{d})$.
Figure 4: System architecture of the proposed EPD Disaggregated Inference.
@@ 3.2 System Design and Optimization
Figure 4 illustrates the architecture of the EPD Disaggregated inference system. Each pipeline stage (Encoding, Prefill, and Decoding) has independent instances that run the corresponding stage. These instances operate in data parallel (DP) mode, enabling concurrent processing of multiple requests per stage and ensuring scalability and efficiency.
@@
In the encoding stage, workers load only the encoder weights and initialize the Multimodal (MM) cache. In the prefill stage, the LLM weights are loaded, and both the MM and KV caches are required to efficiently manage all the data associated with the request. In the decoding stage, workers load the LLM weights for decoding tasks and use the KV cache.
Cache transfers occur asynchronously when an instance pulls a request from its queue, ensuring the transfer occurs once the downstream instance is ready. Further, our system incorporates several techniques to optimize the performance of the disaggregated LMM pipeline, with a primary focus on ensuring smooth token transfers between stages (Section 3.2.1), reducing latency (Section 3.2.2), managing resources effectively (Section 3.2.3), and adapting to changing workloads (Section 3.2.4). The ablation of these features are shown in Sec. 4.4.
@@ 3.2.1 Asynchronous Token Transfer
Disaggregating the system introduces an additional step of transferring vision tokens from the encoding to the prefill stage. To minimize latency during token transfers between stages, our system employs direct, asynchronous transfers via high-bandwidth channels (NVLink, and InfiniBand). The asynchronous transfer allows the system to continue processing new requests without interruption. Both encoding and prefill workers maintain an MM cache to facilitate this process.
When encoding is complete, tokens are stored in the encoding worker’s MM cache, allowing it to serve new requests immediately. An asynchronous event loop monitors completed encoding tasks and initiates direct token transfers to the prefill worker’s MM cache. Once the transfer is confirmed, the encoding cache entries are cleared to free memory. To manage these cache blocks effectively, we introduce the MMBlockManager, which pre-allocates cache blocks based on each request’s needs. After token transfer, the blocks are reassigned or de-allocated, ensuring flexible cache utilization even under heavy workloads.
@@ 3.2.2 Intra-Request Parallel (IRP)
Multimodal requests often include multiple high-resolution images in many practical scenarions like autonomous driving and video question answering. In modern LMMs, these images are further converted into a large number of patches that needs to be processed through a computationally-heavy MME. This significantly increases the computation load, making the encoding process a major bottleneck.
To address this, we introduce Intra-Request Parallelism (IRP), which partitions a single request’s image patches across multiple encoding workers in a data-parallel fashion. Since patches are encoded independently, they can be processed and transferred concurrently. Specifically, each encoding worker concurrently processes a subset of patches, computes their token representations, and asynchronously transfers them to the prefill stage. Once all patch-level tokens reach the prefill stage, they are aligned, projected, and merged to form the complete multimodal tokens.